OpenStudy (anonymous):

A circular wire of radius 3 cm is cut and bent so as to lie along the circumference of hoop with radius 48 cm . find in degrees the angle which is subtended at the center of the hoop?

OpenStudy (aravindg):

what did u try in this question shameer?

OpenStudy (ghazi):

\[\theta= \frac{ arc }{ radius }=\frac{ 3 }{ 48 }=\frac{ 1 }{ 16 }\] i hope you can do the rest :) sorry if i am mistaken

OpenStudy (ghazi):

convert this in to radians ...or whatever you like :)

OpenStudy (anonymous):

Circumfrence = 2 pie * r = 6 pie theta = L/r 6 pie / 48 = pie / 8

OpenStudy (anonymous):

@ghazi i think..u made a mistake...

OpenStudy (anonymous):

@Yahoo! here theta= arc/ radius , where arc = 2*pi*3=6.24*3=18.72

OpenStudy (anonymous):

It is 30 degree right?

OpenStudy (anonymous):